OpenStudy (ajanijones):

What is B Angles A and B are supplementary. m∠A = 24° m∠B =

OpenStudy (ajanijones):

help i give medals

OpenStudy (anonymous):

Supplementary angles are angles that add up to 180 degrees.

OpenStudy (anonymous):

If angle A is 24 degrees and they add up to 180, it's simple subtraction.

OpenStudy (ajanijones):

i know that. I just need to check my answer

OpenStudy (anonymous):

Supplementary angles add up to 180degrees. We can say A+B = 180 So 24+B=180 B = 180-24 B = 156degrees

OpenStudy (agent0smith):

If you need to check your answer, it'd help if you... posted your answer.